# Learning Linear Utility Functions From Pairwise Comparison Queries

### v1

@@ 5.2 Estimating Utility Parameters
Because for $x=\frac{\epsilon}{\sqrt{m-1}}\in(0,1),\exp(x)\geq 1+x$, so $\exp(\frac{-\epsilon}{\sqrt{m-1}})\leq\frac{1}{1+\frac{\epsilon}{\sqrt{m-1}}}$. And $\mathcal{O}(\exp(\frac{-\epsilon}{\sqrt{m-1}})\log(\frac{1}{\delta}))$ is in $\mathcal{O}((\frac{\sqrt{m}}{\epsilon})\log(\frac{1}{\delta})).$
Comparing this result with the passive learning result from inequality (3), whose sample complexity is $\mathcal{O}(\frac{1}{\epsilon}\log(\frac{1}{\delta})+\frac{m}{\epsilon})$, our sample complexity bound is much better since for any $m\in\mathbb{N}$ and $10<\epsilon<$ 1, $\epsilon^{\sqrt{m-1}}<\epsilon<1<\exp(\epsilon),$ implying $\frac{1}{\epsilon}\log(\frac{1}{\delta})>\frac{1}{\exp(\frac{\epsilon}{\sqrt{m-1}})}\log(\frac{1}{\delta})$.

### v3

@@ 5.2 Estimating Utility Parameters
Because for $x=\frac{\varepsilon}{\sqrt{m-1}}\in(0,1),\exp(x)\geq 1+x$, so $\exp(\frac{-\varepsilon}{\sqrt{m-1}})\leq\frac{1}{1+\frac{\varepsilon}{\sqrt{m-1}}}$. And $\mathcal{O}(\exp(\frac{-\varepsilon}{\sqrt{m-1}})\log(\frac{1}{\delta}))$ is in $\mathcal{O}((\frac{\sqrt{m}}{\varepsilon})\log(\frac{1}{\delta})).$
Comparing this result with the passive learning result from inequality (3), whose sample complexity is $\mathcal{O}(\frac{1}{\varepsilon}\log(\frac{1}{\delta})+\frac{m}{\varepsilon})$, our sample complexity bound is much better since for any $m\in\mathbb{N}$ and $10<\varepsilon<$ 1, $\varepsilon^{\sqrt{m-1}}<\varepsilon<1<\exp(\varepsilon),$ implying $\frac{1}{\varepsilon}\log(\frac{1}{\delta})>\frac{1}{\exp(\frac{\varepsilon}{\sqrt{m-1}})}\log(\frac{1}{\delta})$.